OpenStudy (anonymous):

X | y. Which equation is 1 | -1. Represented by the 2 | -3. Table? 3 | -5. Y= 1- 2x 4 | -7. Y= -x -1 Y= x - 2 Y= 2x - 1

OpenStudy (anonymous):

Start with the first equation, y = 1 - 2x Plug in each x to this first equation and calculate y. See if you get the y in the table that is next to the x you plugged in. If it works for all the x's, that's your answer. If not, try the next equation.

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

You don't know how to plug numbers into equations?

OpenStudy (anonymous):

No I don't

OpenStudy (anonymous):

Take the first equation: y = 1 - 2x. Plug in 1 for the x. Perform the calculations. What does y equal?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Plug in 1 for x. That means replace the x with 1...like this: y = 1 - 2x y = 1 - 2(1) Now, can you calculate y?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

yes! and that matches the y value in the table next to x=1. great! Now do the same thing for the other x's in the table. I you always get the y next to the x, the first equation is the answer. If not, go on to the next equation. Good luck!